OpenStudy (anonymous):

If x=at^2 and y=2at then dy/dx =? Plz solve

OpenStudy (agent47):

dx/dt=2at dy/dt=2a dy/dx=(dy/dt)/(dx/dt)=2a/(2at)=1/t

OpenStudy (agent47):

dy/dx = 2a/2at=1/t

OpenStudy (anonymous):

Firstly calculate for dx/dt = 2at ---------1 And dy/dt = 2a ----------------2 Just divide 2 by 1: dy/dx = 2a/2at dy/dx = 1/t is the answer...